A saturated solution of o-nitrophenol has PH=4.53. Calculate its solubility. PKa=7.23

Grade:12

1 Answers

Ramesh V
14 years ago

C6H5NO3----> C6H4NO3- + H+

                                    x                       x

lets solubility be x

so PH = 4.53 or [H+] = 10-4.53  so x = 10-4.53

PKa = 7.23 or [H+] = 10-7.23

Ka = [C6H4NO3-] [H+]   

[C6H4NO3-] = Ka / [H+]
                           = 10 -7.23+4.53

                        solubility   = 2x10-3 g/L